OpenStudy (anonymous):

\[\sqrt{-12} = \sqrt{4\times3\times-1}\]\[\sqrt{4\times \sqrt{3\times \sqrt{-1}}}\]\[2\times \sqrt{3\times \sqrt{-1}}\]\[2i \sqrt{3}\]

OpenStudy (anonymous):

is that the answer ?

OpenStudy (anonymous):

yes the answer is right but the work is wrong

OpenStudy (anonymous):

how do i fix it?

OpenStudy (anonymous):

\[\sqrt{-12}=\sqrt{12}\times \sqrt{-1}=\sqrt{12}i\]

OpenStudy (anonymous):

and \[sqrt{12}=\sqrt{4\times 3}=\sqrt{4}\times \sqrt{3}=2\sqrt{3}\]

OpenStudy (anonymous):

it may be just a typo on your part, but you wrote \[\sqrt{4\times \sqrt{3\times \sqrt{-1}}}\]

OpenStudy (anonymous):

with nested radicals. maybe that is not what you meant because your answer is correct

OpenStudy (anonymous):

the thing wouldn't let put the little square roots, im new to this.

OpenStudy (anonymous):

ok i thought as much. you have it correct if you write it with pencil and paper i am sure

OpenStudy (anonymous):

okay, thank you(: